# Cell-cell communication inference and analysis: biological mechanisms, computational approaches, and future opportunities

### v1

@@ 3 Modeling Strategies for Cell-Cell Communication
Figure 3: General strategies and computational methodologies for inferring CCC from gene expression data. (a) A typical workflow for CCC inference. Gene expression profiles from single-cell or spatial transcriptomics serve as proxies for protein abundance. These data are integrated with prior knowledge, including genes encoding interacting ligands and receptors, co-factors, transcription factors and target genes, to infer CCC. A scoring function quantifies the likelihood or strength of potential CCC events, followed by statistical testing to assess significance. The final output predicts CCC at single-cell or cell-group resolution. (b) Common computational methodologies. The five primary strategies for CCC inference include statistical methods, network methods, deep learning, optimal transport, and factorization methods. Created in https://BioRender.com
Moreover, sophisticated computational methods can be developed by considering diverse biological mechanisms and experimental designs. First, many ligand-receptor binding operates in a multi-subunit architecture, and therefore their heteromeric complexes should be accurately represented. Second, cofactors, including soluble agonists and antagonists, and co-stimulatory and co-inhibitory membrane-bound receptors, introduce additional effects on the core interaction between ligands and receptors, which can be incorporated into CCC modeling. Indeed, many signaling pathways, such as BMP, WNT and TGF beta, are prominently modulated by their cofactors, both positively and negatively[heldin2016signals]. Third, downstream signaling response is often key indicators of whether a cell has genuinely received a signal, greatly helping to mitigate false positives and establish causal relationships in signal transduction. Fourth, since secreting signaling and contact-dependent signaling operate over different spatial ranges, spatial distance should also be taken into account. Finally, comparative analysis across complex experimental designs—including multiple biological replicates and conditions—is essential. For example, analyzing temporal data can reveal the dynamic evolution of CCC, while cross-condition comparisons enable the systematic identification of altered signaling mechanisms and potential therapeutic targets.
@@ 4 Computational Methods for Inferring Cell-Cell Communication
The advent of single-cell and spatially resolved omics technologies, especially transcriptomics, has driven an abundance of computational tools for modeling CCC (Figure 4). These tools exhibit growing methodological diversity and vary in the biological features they are designed to investigate (Figure 5). This diversity reflects ongoing efforts and innovations by researchers to explore CCC mechanisms from multiple perspectives and to develop analytical approaches tailored to specific biological questions. Existing methods for CCC analysis can be broadly categorized into five groups based on their underlying inference strategies (Figure 3b): (1) statistical methods, (2) network methods, (3) deep learning methods, (4) optimal transport methods, and (5) factorization methods. For each category, we first outline the core modeling principles and then evaluate how existing tools address key biological questions. This evaluation is structured around five analytical aspects: spatial constraints, single-cell resolution, intracellular signaling, temporal dynamics, and cross-condition comparison.
@@ 4.1 Statistical Methods
Tools implementing the first strategy primarily diverge in their design of CCC scores and statistical tests. Some apply explicit rules to constrain signaling gene expression, producing binary interaction scores[RN332, RN334, RN370]. More commonly, tools generate continuous scores by calculating the mean, product, or correlation of ligand and receptor expression[RN263, RN338, RN277, RN343, RN266, RN313, RN344, RN289, zheng2022mebocost, jakobsson2021scconnect, RN380, RN414, shao2025extracellular, burdziak2023epigenetic]. The significance of these scores is typically assessed against a null distribution by randomly permuting cell labels. Specific scoring functions reflect distinct biological considerations. CellPhoneDB[RN343], for instance, defines a communication score as the mean of the average ligand expression in one cluster and the average receptor expression in another. To account for protein complexes, it uses the minimum average expression among subunits of ligands or receptors. Other tools incorporate more refined biological models. Given the additional modulation of co-factors (e.g., agonists, antagonists, co-receptors) on the core interaction between ligands and receptors, CellChat[RN342] estimates the level of multimeric ligands and receptors by the geometric mean of their subunits, and designs its scoring function based on the law of mass action to integrate all known molecular interactions. Specifically, the communication probability $P_{i,j}$ from cell groups $i$ to $j$ for a ligand-receptor pair $k$ is modeled as:
@@
Figure 5: Methodological phylogeny of CCC tools. The evolutionary tree classifies 143 computational methods by their core computational strategies (main branches, colored) and the specific biological questions they address (sub-branches). This visualization, inspired by a previous review[RN447], illustrates the evolutionary relationships and functional diversity of CCC tools.
@@
The second strategy employs more advanced statistical methods to model CCC[RN354, RN355, RN414, RN275BulkSignalR, RN356cell2cell, RN367MESSI, RN386]. Several tools within this category estimate communication strength through formal statistical frameworks. For instance, CCPLS[RN355] applies partial least squares (PLS) regression, interpreting regression coefficients as the influence of neighboring cell types—mediated by CCC—on highly variable gene expression. FastCCC[RN414] offers a highly scalable, permutation-free approach that uses fast Fourier transformation (FFT)-based convolution for rapid p-value calculation and a modular scoring system. Other tools utilize statistical methods primarily for the interpretation of inferred interactions. For example, after evaluating the significance of LRIs using null distributions of Spearman correlation coefficients between ligands and receptors, as well as between receptors and their downstream target genes, BulkSignalR[RN275BulkSignalR] associates LRIs with specific cell types by constructing a LASSO regression model. Building upon communication score computation, cell2cell[RN356cell2cell] employs a genetic algorithm to prioritize ligand-receptor pairs that best correlate with physical distance. Additionally, MESSI[RN367MESSI] identifies both intracellular and intercellular signaling genes using a Mixture of Experts (MoE) model with multi-task learning. Topic Modeling[RN386] infers genes affected by intercellular interactions through latent Dirichlet allocation (LDA). Collectively, these tools highlight the diversity of statistical approaches used in CCC analysis.
@@
Spatial omics provides unprecedented opportunity to systematically investigate CCC on native tissues. Incorporating spatial context into CCC analysis allows the calculation of communication scores to be constrained within a short distance range, aligning with the intuitive assumption that cells in close proximity are more likely to interact. Giotto[RN364Giotto] identifies significantly interacting cell type pairs by randomly permuting cell type labels within a predefined spatial neighborhood network. Communication scores are then calculated only for cell pairs deemed spatially proximal and likely to interact based on cell type. CellNeighborEX[RN358CellNeighborEX] leverages spatial transcriptomics data to identify neighbor-dependent gene expression by comparing transcriptomes in heterotypic versus homotypic cellular contexts, providing an unbiased approach to uncover genes regulated by direct cell contact and the immediate microenvironment, thereby revealing a layer of intercellular communication that complements and extends beyond canonical ligand-receptor analysis. SpaCET[RN374SpaCET] deconvolves immune and stromal cell lineage fractions in tumors by integrating cancer-specific genomic patterns and hierarchical regression, and it infers CCC by requiring both spatial colocalization and significant ligand-receptor co-expression within spots, providing a spatially grounded framework to analyze interactions in the tumor microenvironment. Niche-DE[RN423Niche-DE] is a regression-based statistical framework that identifies cell-type-specific genes whose expression is modulated by the local cellular context, and reveals LRIs that drive niche-differential expression patterns from spatial transcriptomics data.
@@
Most tools compute communication scores using gene expression profiles aggregated at the cell group level. Although this strategy helps mitigate the sparsity of scRNA-seq data, it fails to fully exploit the advantages of single-cell resolution. Consequently, it cannot capture cell-specific interactions within cell groups and may introduce bias due to predefined cell clusters. To address these limitations, several recent methods have been developed to infer interactions directly between individual cell (or spot) pairs[RN303, RN410CytoSignal, RN432Spacia]. stLearn[RN382stLearn] integrates spatial information to compute communication scores in both ‘within-spot’ and ‘between-spot’ modes, while also accounting for cell type diversity to facilitate analyses across diverse biological contexts. Spacia[RN432Spacia] directly models CCC at single-cell resolution using a multi-instance learning framework and explicitly captures complex many-to-one relationships between sender and receiver cells. Conceptually different from other CCC inference methods, SpatialDM[RN378SpatialDM] explicitly incorporates spatial distance through a spatial weight matrix. It identifies significant ligand-receptor pairs and their local interaction hotspots using a bivariate Moran’s I statistic, bypassing the need for permutation tests via an analytical null distribution. This makes SpatialDM both statistically rigorous and highly scalable to datasets comprising millions of cells. CytoSignal[RN410CytoSignal] predicts signaling locations and activity at single-cell resolution by computing a weighted average of ligand expression across a cell’s neighborhood and then multiplying by the receptor expression in the target cell.
@@
Based on the established importance of analyzing CCC, understanding the subsequent intracellular signaling cascades and transcriptional regulation is crucial for deciphering the complete cellular response mechanism. To bridge this gap, computational methods increasingly leverage curated pathway databases, such as KEGG[kanehisa2000kegg] and Reactome[croft2010reactome], to connect LRIs to potential downstream genes through hypothesis testing or correlation analysis[RN282, RN312, RN286, RN345]. Among these tools, SingleCellSignalR[RN286] infers CCC from scRNA-seq data using a curated ligand-receptor database and a regularized score that provides a stable threshold to control false positives, facilitating the identification of high-confidence interactions for downstream validation. A more integrated approach is exemplified by CellPhoneDB v5[RN316CellPhoneDBv5]. It extends beyond LRI identification with its CellSign module[RN345], which is designed to connect interactions to intracellular signaling and transcriptional outcomes. This module utilizes curated knowledge of receptor-TF relationships and downstream TF activity data to model signal propagation. DcjComm[RN411] is a versatile computational tool that employs an NMF-based joint learning model to simultaneously identify functional gene modules, perform dimension reduction, and cluster cells, followed by a statistical model to infer CCC by integrating ligand-receptor pairs with downstream TF-target gene activities. Other tools incorporate the expression of downstream genes directly into the calculation of communication scores to provide a more holistic inference of CCC, such as SoptSC[wang2019cell], CellCall[RN268], BulkSignalR[RN275BulkSignalR] and IGAN[RN417IGAN]. The recent method BulkSignalR[RN275BulkSignalR] addresses the challenge of inferring CCC from bulk and spatial multi-cellular data by statistically integrating LRIs with downstream pathway activity, providing a robust and functionally contextualized approach to decipher cellular networks from these prevalent data types. IGAN[RN417IGAN] leverages spatial transcriptomics to construct intercellular gene association networks between adjacent cells by performing a non-parametric statistical model. It further elucidates the full upstream and downstream pathway context and thereby reveals communication heterogeneity and pathway-level mechanisms.
@@
Incorporating temporal dynamics into CCC analysis enables a more nuanced understanding of cell state transitions and extends the analytical scope of computational tools, which is particularly valuable in contexts such as embryonic development, cancer progression, and other temporally evolving biological processes. stLearn[RN382stLearn] introduces an framework that integrates gene expression with spatial information to reconstruct spatiotemporal trajectories, thereby uncovering spatial branching events that occur over time within tissues. Beyond trajectory-focused tools[RN382stLearn, wang2019cell], TimeTalk[RN282] combines trajectory inference with causal testing to identify regulatory relationships between ligand-receptor pairs and temporally relevant TFs during early embryogenesis. DIISCO[RN413], on the other hand, utilizes Gaussian process regression models to characterize dynamic changes in cell type proportions and to predict time-resolved intercellular communication patterns.
@@
Advances in experimental technologies have increased the availability of multi-condition datasets, spurring demand for computational methods that compare CCC across biological contexts. Such comparative analysis is crucial for identifying altered signaling mechanisms that drive cell fate decisions and phenotypic transitions. Common strategies for this task involve coupling differential CCC network analysis with differential gene expression profiling[RN332, RN301, zheng2022mebocost, zhao2023inferring], or applying statistical tests (e.g., Fisher’s exact test, permutation tests) to communication scores[RN334, RN263, RN338, RN407]. Beyond these approaches, CellChat[RN342, jin2025cellchat] employs network topology analysis, using centrality measures to identify key signaling sources, targets, mediators, and influencers within a network, and further groups pathways based on functional and topological similarity. Several tools are specifically designed for differential CCC analysis. scDiffCom[RN263] provides a statistically rigorous framework to identify dysregulated LRIs and cell-type pairs between conditions (e.g., aging or disease), powering the scAgeCom atlas of age-related communication changes in mouse tissues. Community[RN407] uniquely integrates cell type abundance, active fraction, and expression level to pinpoint altered interactions and disentangle compensatory mechanisms in case-control studies. CellCallEXT[gao2022cellcallext] extends its precursor CellCall[RN268] by incorporating transcription factor activity alterations between conditions to infer disease-associated changes in both intercellular communication and intracellular signaling. To address scalability in large-scale studies, FastCCC[RN414] introduces a permutation-free framework for analytical p-value computation. Its key innovation is a reference-based paradigm that leverages a large human atlas, providing a healthy baseline to robustly identify dysregulated communications in disease. CACIMAR[RN403] proposes a conservation score to identify CCC events that are conserved across different species. For spatial transcriptomics, SpatialDM[RN378SpatialDM] uses generalized linear models to test if experimental conditions significantly influence interaction density across multiple samples.

### v3

@@ 3 Modeling Strategies for Cell-Cell Communication
Figure 3: General strategies and computational methodologies for inferring CCC from gene expression data. (a) A typical workflow for CCC inference. Gene expression profiles from single-cell or spatial transcriptomics serve as proxies for protein abundance. These data are integrated with prior knowledge, including genes encoding interacting ligands and receptors, co-factors, transcription factors and target genes, to infer CCC. A scoring function quantifies the likelihood or strength of potential CCC events, followed by statistical testing to assess significance. The final output predicts CCC at single-cell or cell-group resolution. (b) Common computational methodologies. The five primary strategies for CCC inference include statistical methods, network methods, deep learning, optimal transport, and factorization methods. Created in BioRender. S. Jin, 2026, https://BioRender.com/llyvm5g.
Moreover, sophisticated computational methods can be developed by considering diverse biological mechanisms and experimental designs. First, many ligand-receptor interactions operate in a multi-subunit architecture, and therefore their heteromeric complexes should be accurately represented. Second, cofactors, including soluble agonists and antagonists, and co-stimulatory and co-inhibitory membrane-bound receptors, introduce additional effects on the core interaction between ligands and receptors, which can be incorporated into CCC modeling. Indeed, many signaling pathways, such as BMP, WNT and TGF-$\beta$, are prominently modulated by their cofactors, both positively and negatively[73]. Third, downstream signaling response is often a key indicator of whether a cell has genuinely received a signal, greatly helping to mitigate false positives and establish causal relationships in signal transduction. Fourth, since secreted signaling and contact-dependent signaling operate over different spatial ranges, spatial distance should also be taken into account. Finally, comparative analysis across complex experimental designs—including multiple biological replicates and conditions—is essential. For example, analyzing temporal data can reveal the dynamic evolution of CCC, while cross-condition comparisons enable the systematic identification of altered signaling mechanisms and potential therapeutic targets.
@@ 4 Computational Methods for Inferring Cell-Cell Communication
The advent of single-cell and spatially resolved omics technologies, especially transcriptomics, has driven an abundance of computational tools for modeling CCC (Figure 4). These tools exhibit growing methodological diversity and vary in the biological features they are designed to investigate (Figure 5). This diversity reflects ongoing efforts and innovations by researchers to explore CCC mechanisms from multiple perspectives and to develop analytical approaches tailored to specific biological questions. Existing methods for CCC analysis can be broadly classified into five categories based on their core inference strategies (Figure 3b): (1) statistical methods, (2) network methods, (3) deep learning methods, (4) optimal transport methods, and (5) factorization methods. For each category, we first outline the core modeling principles and then evaluate how current tools address key biological questions. This evaluation is structured around five analytical aspects: spatial constraints, cellular resolution (e.g., single-cell level or spot-level resolution), intracellular signaling, temporal dynamics, and cross-condition comparison.
@@ 4.1 Statistical Methods
Tools implementing the first strategy primarily diverge in their design of CCC scores and statistical tests. Some apply explicit rules to constrain signaling gene expression, producing binary interaction scores[37, 176, 62]. More commonly, tools generate continuous scores by calculating the mean, product, or correlation of ligand and receptor expression[99, 111, 188, 51, 168, 113, 133, 91, 208, 81, 136, 78, 154, 26]. The significance of these scores is typically assessed against a null distribution by randomly permuting cell labels. Specific scoring functions reflect distinct biological considerations. CellPhoneDB[51], for instance, defines a communication score as the mean of the average ligand expression in one cluster and the average receptor expression in another. To account for protein complexes, it uses the minimum average expression among subunits of ligands or receptors. Other tools incorporate more refined biological models. Given the additional modulation of co-factors (e.g., agonists, antagonists, co-receptors) on the core interaction between ligands and receptors, CellChat[88] estimates the level of multimeric ligands and receptors by the geometric mean of their subunits, and designs its scoring function based on the law of mass action to integrate all known molecular interactions. Specifically, the communication probability $P_{i,j}$ from cell groups $i$ to $j$ for a ligand-receptor pair $k$ is modeled as:
@@
Figure 5: Methodological phylogeny of CCC tools. The evolutionary tree classifies 143 computational methods by their core computational strategies (main branches, colored) and the specific biological questions they address (sub-branches). This visualization, inspired by a previous review[9], illustrates the evolutionary relationships and functional diversity of CCC tools.
@@
The second strategy employs more advanced statistical methods to model CCC[185, 165, 78, 172, 11, 102, 137]. Several tools within this category estimate communication strength through formal statistical frameworks. For instance, CCPLS[165] applies partial least squares (PLS) regression, interpreting regression coefficients as the influence of neighboring cell types—mediated by CCC—on highly variable gene expression. FastCCC[78] offers a highly scalable, permutation-free approach that uses fast Fourier transformation (FFT)-based convolution for rapid p-value calculation and a modular scoring system. Other tools utilize statistical methods primarily for the interpretation of inferred interactions. For example, after evaluating the significance of LRIs using null distributions of Spearman correlation coefficients between ligands and receptors, as well as between receptors and their downstream target genes, BulkSignalR[172] associates LRIs with specific cell types by constructing a LASSO regression model. Building upon communication score computation, cell2cell[11] employs a genetic algorithm to prioritize ligand-receptor pairs that best correlate with physical distance. Additionally, MESSI[102] identifies both intracellular and intercellular signaling genes using a Mixture of Experts (MoE) model with multi-task learning. Topic Modeling[137] infers genes affected by intercellular interactions through latent Dirichlet allocation (LDA). Collectively, these tools highlight the diversity of statistical approaches used in CCC analysis.
@@
Spatial omics provides an unprecedented opportunity to systematically investigate CCC on native tissues. Incorporating spatial context into CCC analysis allows the calculation of communication scores to be constrained within a short distance range, aligning with the intuitive assumption that cells in close proximity are more likely to interact. Giotto[50] identifies significantly interacting cell type pairs by randomly permuting cell type labels within a predefined spatial neighborhood network. Communication scores are then calculated only for cell pairs deemed spatially proximal and likely to interact based on cell type. CellNeighborEX[94] leverages spatial transcriptomics data to identify neighbor-dependent gene expression by comparing transcriptomes in heterotypic versus homotypic cellular contexts, providing an unbiased approach to uncover genes regulated by direct cell contact and the immediate microenvironment, thereby revealing a layer of intercellular communication that complements and extends beyond canonical ligand-receptor analysis. SpaCET[148] deconvolves immune and stromal cell lineage fractions in tumors by integrating cancer-specific genomic patterns and hierarchical regression, and it infers CCC by requiring both spatial colocalization and significant ligand-receptor co-expression within spots, providing a spatially grounded framework to analyze interactions in the tumor microenvironment. Niche-DE[122] is a regression-based statistical framework that identifies cell-type-specific genes whose expression is modulated by the local cellular context, and reveals LRIs that drive niche-differential expression patterns from spatial transcriptomics data.
@@
Most tools compute communication scores using gene expression profiles aggregated at the cell group level. Although this strategy helps mitigate the sparsity of scRNA-seq data, it fails to fully exploit the advantages of single-cell resolution. Consequently, it cannot capture cell-specific interactions within cell groups and may introduce bias due to predefined cell clusters. To address these limitations, several recent methods have been developed to infer interactions directly between individual cell (or spot) pairs[147, 110, 210]. stLearn[142] integrates spatial information to compute communication scores in both ‘within-spot’ and ‘between-spot’ modes, while also accounting for cell type diversity to facilitate analyses across diverse biological contexts. Spacia[210] directly models CCC at single-cell resolution using a multi-instance learning framework and explicitly captures complex many-to-one relationships between sender and receiver cells. Conceptually different from other CCC inference methods, SpatialDM[107] explicitly incorporates spatial distance through a spatial weight matrix. It identifies significant ligand-receptor pairs and their local interaction hotspots using a bivariate Moran’s I statistic, bypassing the need for permutation tests via an analytical null distribution. This makes SpatialDM both statistically rigorous and highly scalable to datasets comprising millions of cells. CytoSignal[110] predicts signaling locations and activity at single-cell resolution by computing a weighted average of ligand expression across a cell’s neighborhood and then multiplying by the receptor expression in the target cell.
@@
Based on the established importance of analyzing CCC, understanding the subsequent intracellular signaling cascades and transcriptional regulation is crucial for deciphering the complete cellular response mechanism. To bridge this gap, computational methods increasingly leverage curated pathway databases, such as KEGG[93] and Reactome[40], to connect LRIs to potential downstream genes through hypothesis testing or correlation analysis[177, 190, 27, 61]. Among these tools, SingleCellSignalR[27] infers CCC from scRNA-seq data using a curated ligand-receptor database and a regularized score that provides a stable threshold to control false positives, facilitating the identification of high-confidence interactions for downstream validation. A more integrated approach is exemplified by CellPhoneDB v5[164]. It extends beyond LRI identification with its CellSign module[61], which is designed to connect interactions to intracellular signaling and transcriptional outcomes. This module utilizes curated knowledge of receptor-TF relationships and downstream TF activity data to model signal propagation. DcjComm[48] is a versatile computational tool that employs an NMF-based joint learning model to simultaneously identify functional gene modules, perform dimension reduction, and cluster cells, followed by a statistical model to infer CCC by integrating ligand-receptor pairs with downstream TF-target gene activities. Other tools incorporate the expression of downstream genes directly into the calculation of communication scores to provide a more holistic inference of CCC, such as SoptSC[178], CellCall[201], BulkSignalR[172] and IGAN[211]. The recent method BulkSignalR[172] addresses the challenge of inferring CCC from bulk and spatial multi-cellular data by statistically integrating LRIs with downstream pathway activity, providing a robust and functionally contextualized approach to decipher cellular networks from these prevalent data types. IGAN[211] leverages spatial transcriptomics to construct intercellular gene association networks between adjacent cells by performing a non-parametric statistical model. It further elucidates the full upstream and downstream pathway context and thereby reveals communication heterogeneity and pathway-level mechanisms.
@@
Incorporating temporal dynamics into CCC analysis enables a more nuanced understanding of cell state transitions and extends the analytical scope of computational tools, which is particularly valuable in contexts such as embryonic development, cancer progression, and other temporally evolving biological processes. stLearn[142] introduces a framework that integrates gene expression with spatial information to reconstruct spatiotemporal trajectories, thereby uncovering spatial branching events that occur over time within tissues. Beyond trajectory-focused tools[142, 178], TimeTalk[177] combines trajectory inference with causal testing to identify regulatory relationships between ligand-receptor pairs and temporally relevant TFs during early embryogenesis. DIISCO[184], on the other hand, utilizes Gaussian process regression models to characterize dynamic changes in cell type proportions and to predict time-resolved intercellular communication patterns.
@@
Advances in experimental technologies have increased the availability of multi-condition datasets, spurring demand for computational methods that compare CCC across biological contexts. Such comparative analysis is crucial for identifying altered signaling mechanisms that drive cell fate decisions and phenotypic transitions. Common strategies for this task involve coupling differential CCC network analysis with differential gene expression profiling[37, 34, 208, 206], or applying statistical tests (e.g., Fisher’s exact test, permutation tests) to communication scores[176, 99, 111, 157]. Beyond these approaches, CellChat[88, 89] employs network topology analysis, using centrality measures to identify key signaling sources, targets, mediators, and influencers within a network, and further groups pathways based on functional and topological similarity. Several tools are specifically designed for differential CCC analysis. scDiffCom[99] provides a statistically rigorous framework to identify dysregulated LRIs and cell-type pairs between conditions (e.g., aging or disease), powering the scAgeCom atlas of age-related communication changes in mouse tissues. Community[157] uniquely integrates cell type abundance, active fraction, and expression level to pinpoint altered interactions and disentangle compensatory mechanisms in case-control studies. CellCallEXT[60] extends its precursor CellCall[201] by incorporating transcription factor activity alterations between conditions to infer disease-associated changes in both intercellular communication and intracellular signaling. To address scalability in large-scale studies, FastCCC[78] introduces a permutation-free framework for analytical p-value computation. Its key innovation is a reference-based paradigm that leverages a large human atlas, providing a healthy baseline to robustly identify dysregulated communications in disease. CACIMAR[85] proposes a conservation score to identify CCC events that are conserved across different species. For spatial transcriptomics, SpatialDM[107] uses generalized linear models to test if experimental conditions significantly influence interaction density across multiple samples.